Aggregator

Autolog save

6 days 23 hours ago
Log Number: 4467519Logbook(s): Tags: Body: saver ran from opsl632

Saved: samari 01-28-26 11:52:46 {A:5 B:3 C:1 E:340 A=BS B=BS C=BS D=BS 0R08=0.00 0R08 dump in. Autosave prior to restore.
samari

We are now using JLab HV Drivers

6 days 23 hours ago
Log Number: 4467516Logbook(s): Entry Makers: suleimanBody: 

Figure 1: New RTP Driver - This is the new control screen.
Figure 2: Old RTP Cell - This is what we have been using with the UVA Driver.

Both screens are still available on JMenu from: Injector -> Parity

suleiman

Access of E3, E5, W3, and W5 racks

6 days 23 hours ago
Log Number: 4467515Logbook(s): Body: 

E. Diggs and M. Fratarcangeli entered W3, W5, E3, and E5 to turn on fans for the rack. M. Fratarcangeli unlocked racks W501B06, W301B08, E301B09, and E501B09 for the work to be completed and secured them upon the completion of the work (see attached photos of secured locks post work). SRPDs were worn and read 0.0 at the end of the work.

melissaf

COMMENT ON VME/CAMAC Preventative maintenance (SAM 2025-2026)

6 days 23 hours ago
Log Number: 4467514Logbook(s): Body: 

During my initial walk I apparently wrote in the wrong space on my PM sheet. The other fan tray needing to be swapped is IOCNL2 and NOT IOCNL1B. IOCNL2's crate has VMIC cards which have sensitive connectors and with us being so close to lockup after lunch I will postpone IOCNL2's swap until tomorrow as I'd like to have the owners of these cards do the manipulation required to get the fan tray out. However, It is the same systems involved and I will reach out to the system owners to re-coordinate for a swap tomorrow.

sawyer

COMMENT ON YBBS04 & 06 Hose Repair/Replace

6 days 23 hours ago
Log Number: 4467513Logbook(s): Body: 

After leak repairs are completed, we plan to circulate heated LCW thru the coils to ensure there are no dead headed circuits. This will be achieved using a thermal camara. Water heating will be done using our small flushing rig.

wilsonn

iocnl6mag restored

6 days 23 hours ago
Log Number: 4467512Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-01-28 11:02:25 The output from the myRestore execution was as follows: iocnl6mag magnet_ioc magnet_MAG@2026-01-28 11:02 178(178) restored Pause: 2 second(s) magnet_HVAL@2026-01-28 11:02 178(178) restored Pause: 2 second(s) magnet_LOOP@2026-01-28 11:02 178(178) restored Pause: 2 second(s) magnet_CMPQ@2026-01-28 11:02 372(372) restored magnet_TrimRack@2026-01-28 11:02 12(12) restored Procedure return status: Success systemEnum@2026-01-28 11:02 2(2) restor
alarms

HYST AREA 7-0 Cycles Magnets

6 days 23 hours ago
Log Number: 4467511Logbook(s): Body: 

HYST AREA Version 7-0 has successfully cycled the following:

MQA5C01
MBD5C01H
MBD5C01V
MQA5C02
MQA5C02A
MBD5C02H
MBD5C02V
MQA5C03
MBD5C03H
MBD5C03V
MQA5C04
MBD5C04H
MBD5C04V
MBD5C05H
MBD5C05V
MBD5C06H
MBD5C06V
MBD5C07H
MBD5C07V
MBD5C08H
MBD5C08V
MBD5C00H
MBD5C00V

Total Time: 3.050 mins

banks

Shift Information

Shift Day
CEBAF Program
MCC Crew Chief Humphry
MCC Operators Janak
Program Deputy Mock
Announcements

Access Status

Recent PSS History

recent pss history graph